Circular No. 8902

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 W4 = P/1993 D1 (HILL) S. Foglia, R. Matson, and M. Tombelli report the identification of “pre-

covery” images of comet P/2006 W4 (cf. IAUC 8779), aided by the orbital elements published on MPEC 2007-X14 (with indicated ΔT = +1.06 days when run back to 1992), and their astrometric measurements are provided below from the trails on the two U.K. Schmidt Telescope plates.

1993 UT  α_{2000}  δ_{2000}
Feb. 26.53282  9^{h}44^{m}59.^{s}08  −34^{o}38′11′′.0
26.57795  9 44 57.60  −34 38 01.8
Apr. 5.58208  13 04 38.14  −37 52 27.4
5.62375  13 04 36.85  −37 52 15.9

The following orbital elements by B. G. Marsden are from 92 observations, 1993–2007 (mean residual 0′′.6):

Epoch = 1992 June 27.0 TT

T = 1992 June 21.6666 TT  ω = 247.^{o}5818
e = 0.314858  Ω = 243.4600 }2000.0
q = 4.416239 AU  i = 36.2938

a = 6.445725 AU n^{o} = 0.0602277 P = 16.36 years Epoch = 2009 Jan. 9.0 TT

T = 2009 Jan. 21.0916 TT  ω = 249.^{o}6184
e = 0.314866  Ω = 243.2496 }2000.0
q = 4.438583 AU  i = 36.3618

a = 6.478413 AU n^{o} = 0.0597725 P = 16.49 years

V597 PUPPIS R. J. Rudy, R. W. Russell, and D. K. Lynch, The Aerospace Corpora-

tion; and C. E. Woodward, University of Minnesota, report 0.8- to 2.5-μm spectroscopy of this nova (cf. IAUC 8895, 8896) using the Infrared Telescope Facility (+ SpeX) on Nov. 30.54 UT. V597 Pup was in its early stage of development. He I 1.0830-μm was very strong, as was He I 20581-μm. The O I lines were strong, and Brγ showed a doubled line profile. There was no evidence of He II lines or thermal emission from dust. The FWHM of the lines was ≈ 3900 km/s.

Visual magnitude estimates by A. Amorim, Florianopolis, Brazil: Nov. 16.047 UT, 8.1; 20.041, 9.7; 22.078, 9.8; 23.040, 10.0.

2007 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green